# Judging the Judges: Evaluating Alignment and Vulnerabilities in LLMs-as-Judges

### v5

@@ 4 Results
Figure 2: Difference with human evaluation scores versus alignment metric. The delta evaluation score is the difference between the judge and the human score; y-axes are in log scale. Percent alignment (left) shows a very skewed distribution, making it difficult to distinguish models. Scott’s $\mathbf{\pi}$ (right) provides a clearer difference between models, and is more indicative of deviation of the gold score.
In this section we discuss our main results, primarily focusing on the relationship between evaluations by various judge models and human evaluations (§ 4.1), and how that impacts their usability (§ 4.2). To do so, we evaluate their alignment with human judgment and assess how differently they rank the nine exam-taker models compared to humans. In Section 5, we further analyse their precision and recall to further investigate the types of errors that can be made by various judge models. Details about compute requirements and others costs for experiments are given in Appendix H.
@@ 4.1 Alignment between judge models and humans
We start by computing Scott’s $\mathbf{\pi}$ scores and percent agreement between the evaluations of each judge model and the human annotators. We show the result in Figure 1. We observe that percent alignment is high for virtually all models, with the exception of Gemma 2B and EM. Scott’s $\mathbf{\pi}$, on the other hand, has low values for most models, though its value is in the high 80s for Llama-3 70B, Llama-3.1 70B and GPT-4 Turbo. Nevertheless, there still is a significant disparity between human judgment and judge models: the best scoring judge, Llama-3 70B, is 8 points behind human judgment. Notably, EM has the most variance in alignment, while Gemma 2B has the lowest alignment amongst all judges.
@@
To understand how indicative the two alignment metrics are of the expected accuracy of the overall judgement of the models, we plot, for each judge model and exam-taker model, the difference between the score assigned by the judge and the score assigned by a human. In the figure, we can see that for Scott’s $\mathbf{\pi}$ values higher than 80, the evaluation scores are comparatively close to the human evaluation scores, with a difference of up to 5 points in their assigned scores (complete results table provided in Appendix J). For percent alignment, on the other hand, even judges that have more than 90% may still differ more than 10 points in their assigned score. Interestingly, the deviation from human-judgements for a single judge model can be quite different depending on the exam-taker model. In Figure 1(a), Gemma 2B, for instance, sometimes assigns higher scores than humans, and sometimes much lower. In the next section, we further explore this particular pattern.
@@ 4.2 Exploring consistent patterns in judge models
In the previous section, we have seen that none of the judge models we considered were aligned with humans as well as the humans were aligned amongst themselves. Furthermore, as can be seen in Figure 2, the scores assigned by even the best aligned judge models can differ up to 5 points with the human-assigned scores. However, while this may limit – to some extent – the utility of using a judge models to get a perfect estimate of the exam-taker model’s capability on the benchmark, the judge models may still offer valuable insights to differentiate between different exam-taker models. If judges exhibit consistent biases such as – akin to a very strict teacher – consistently rating any exam-taker model lower, they will not assign identical scores but may assign identical rankings.
To evaluate this, we compare the rankings assigned by each judge model to the nine exam-taker models by computing their Spearman’s rank correlation coefficients $\rho$ (Spearman, 1904) with the human ranking. We show the rankings in Figure 3(a), with $\rho$ and corresponding $\sigma$ values in Appendix L. Most judge models have rank correlations higher than 0.7; it appears they struggle to distinguish between poorer-performing exam-taker models, but do well at distinguishing between better-performing ones. Notably, the results show that several models that assign scores quite divergent from humans and have poor alignment on the sample level are very aligned in terms of the rankings they assign. Specifically, both contains and Mistral 7B, with Scott’s $\mathbf{\pi}$ values of 64 and 66, respectively, exhibit very high rank correlation with the human scores ($\rho$ 0.99 and 0.98, respectively, with $\sigma$ 0.02 and 0.03). With that, these judges perform on par with GPT-4 Turbo and outperform the better Llama judges – though with lower significance values – indicating that identifying which models are better should not be equated to assigning them the correct score.
Figure 3: Judge rankings and true/false positives and negatives. (a) Assigned exam-taker model rankings assigned by highly human aligned judges. Contains stays closely to human-assigned rankings, as well as GPT-4 Turbo and Mistral 7B. (b) False positives and negatives across different judge models, in descending order of human alignment. Both false negatives and false positives increase as human alignment decreases, but well-aligned models tend to produce more false negatives than false positives.
@@ 5 Analysis
To better understand the judge models, we conduct multiple case studies aimed at identifying common errors and vulnerabilities in the judges we investigate. Specifically, we study their precision and recall and error types (§ 5.1), their sensitivity to the instruction prompt prompt (§ 5.2), how they respond to controlled resposes of specific types (§ 5.3), and the extent to which they have a leniency bias (§ 5.4).
We first investigate the precision and recall of the judge models. Maintaining the ordering of Figure 1, we plot both in Figure 4(a). We can see that both precision and recall exhibit a moderate increasing trend as alignment increases. In Figure 3(b), we observe a similar pattern, though with a clearer picture on the distribution of false positives and negatives. Specifically, we see that the number of true positives is quite stable across many judges. The true negatives, instead, drop off quickly as the judge quality decreases, suggesting it is generally easier to judge answers that are correct.
@@ 5.1 Better aligned models: Precision and recall gains with error spotlights
Next, we analyse the types of errors made by the judge models by manually annotating 900 outputs from Llama-7B Base with error codes, focusing on the top performers GPT-4 Turbo and Llama-3 70B. We then determine the percentage of each error type that are correctly judged to be incorrect by these two models. The results are shown in Table 2, where it can be observed that both GPT-4 Turbo and Llama-3 70B have a good error recall when the answers refer to an incorrect entity, or when too many entities are present. Under-specified and otherwise incorrect answers are most challenging for both judges, while answers with too few entities are judged relatively accurately by GPT-4 but less accurately by Llama-3 70B.
@@ 5.2 Judge model sensitivity to prompt length and specificity
Next, we study the impact of the prompt on the predictions of the judge models, to understand if the success of various judge models is related to the length of the prompt, and to study the degree to which the judgments of the judge models change with the specificity of the prompt. We use four different prompt versions, varying in length and specificity. The first two prompts (Without guidelines V1/V2, 45 and 58 tokens, respectively) simply ask to evaluate the responses, without any further information, while more elaborate guidance and examples are given in the longer prompts (Guidelines without examples and Guidelines with examples, 245 and 301 tokens, respectively). All prompts are listed in Appendix M.
Figure 4(b) shows that GPT-4 Turbo, Llama-3 70B and Llama-3.1 70B exhibit relatively low variance in their agreement with humans as the level of information and the length of the prompt increases. For this task, top performers’ (GPT-4 Turbo, Llama-3 70B and Llama-3.1 70B) implicit definition of a correct judgment seems well aligned with the provided instructions and thus shows high alignment with humans even if no specific instructions are provided. It can also be observed that only top performers appears to benefit from the more detailed instructions, with a slight upward trend, whereas the other models get less aligned with more instructions. This might be due to the less powerful judges not being able to follow many instructions in the prompt at the same time. In a follow-up experiment, we further investigate the impact of the order of the reference answers (for details, we refer to Appendix N). Figure 6(b) illustrates that larger judge models consistently maintain their judgments regardless of the reference order, whereas smaller models – with the exception of Mistral 7B – are more sensitive to the reference order given in the prompt.
@@ 5.3 Evaluating controlled responses
Next, we perform simple tests on the judge models by asking them to evaluate a set of dummy benchmark responses. For the first test, the answer to be evaluated for each question is one of the references from the dataset, verbatim (the answer is thus always correct). For the next three tests, the answer is always incorrect. In the second and third tests, the dummy exam-taker model always responds with ‘‘Yes’’, and ‘‘Sure’’ for the second and third tests, respectively. For the fourth test, the evaluated answer is a repetition of the question.
In Figure 5, we can see that while some judge models are able to identify and correctly mark the answers as correct (for the first test) or incorrect (for the next three tests), some judges, notably Llama-2 70B, incorrectly evaluate a significant number of dummy answers, even though they show a relatively high alignment with humans on the benchmark evaluations (see Figure 1(b)). We hypothesise that when the answers are plausible but incorrect (e.g. if the question asks about the name of the author of a book, and the exam-taker model gives the name of the wrong author), most judges are able to identify them as being incorrect (by comparing it with the reference answer). However, the judges might get confused about what they are supposed to evaluate if the answer is completely unrelated to the question (such as the words ‘‘Yes’’ and ‘‘Sure’’). It is possible that, in this situation, a judge model tries to evaluate one of the reference answers, thus marking it as correct, though further research is required to identify the cause of this behavior.
@@ 5.4 Leniency bias in judge models
Judge model $\pi$ $P_{c}$ $P_{+}$
JudgeLM-7B 0.65 0.68 0.19
Mistral-7B 0.66 0.70 0.87
Llama2-70B 0.69 0.66 0.99
Llama2-13B 0.74 0.74 0.87
Llama3.1-8B 0.77 0.77 0.82
GPT-4 0.87 0.87 0.69
Llama3.1-70B 0.88 0.88 0.82
Llama3-70B 0.88 0.87 0.90
Figure 6: Leniency bias and answer consistency. (a) Leniency bias across various judge models. (b) Consistency score, defined as the percentage of questions for which the judge model gives the same judgment for three different answer orders.
Lastly, to get a general sense of the inherent biases or misalignment in the evaluation criteria that might be present in the judge models, we estimate if they have a positive or negative bias in their judgment. To do so, we assume that a judge assigns the correct judgment (i.e. same evaluation as the ground truth) with a probability of $P_{c}$ and assigns the rest of the samples to be ‘‘correct’’ with a probability $P_{+}$, which we call their leniency bias. We estimate the values of $P_{c}$ and $P_{+}$ from the benchmark results555 The theoretical derivation of the expressions for $P_{c}$ and $P_{+}$, as well as the empirical validation for their estimated values can be found in Appendix O. and show them in Figure 6(a). We observe that $P_{+}$ for most models is significantly higher than $0.5$, indicating a tendency of the judge models to evaluate responses as ‘‘correct’’ when their evaluation criteria are not completely aligned with the provided instructions.
@@ 6 Conclusion
In this work, we provide an extensive study of the properties of LLMs as judges, comparing them with human judges as well as automated evaluation methods. By focusing on a clean evaluation scenario in which inter-human agreement is high, we examine the potential issues with the LLM-as-a-judge paradigm, separately from the ambiguity and subjectivity in the task itself. We find that even in straightforward setups, smaller and more cost-efficient models are less effective judges compared to the best available LLMs, such as Mistral 7B. GPT-4 Turbo, Llama-3.1 70B and Llama-3  70B, instead, are much better aligned, though they are still quite far from the alignment that humans have among each other. In some cases, despite their high alignment, their scores deviate from human scores with up to 5 points. Given the relative simplicity of the scenarios in which we deployed the judges, urging caution in using judge for more complex scenarios. Importantly, we noted that such patterns are virtually undetectable using the commonly deployed metric of percent aligned, which barely discrimates between the considered judges. We suggest that future work instead considers the more robust metric Scott’s $\mathbf{\pi}$, which allows to distinguish judges much better.
@@
Next, we note that to discriminate between models, high alignment scores are not an absolute necessity. While GPT-4 Turbo and Llama-3 both have excellent alignment scores, simpler and more cost-efficient and even the lexical matching method contains perform on par when discriminating between the exam-taker models in terms of their ranking, despite having much lower alignment scores and score deviations. If the purpose of a study is to determine which model is better and not to estimate their actual scores, such approaches may thus be as suitable as the more expensive ones.
@@
Lastly, we run a range of experiments to investigate judge models’ sensitivity to prompts, their precision and recall, their error types, how lenient they are, and how much they can be fooled by dummy answers. We find that LLMs tend to judge positively when in doubt, and this is more pronounced for small models than for larger ones; that judge models with lower alignment lack precision rather than recall, that better models are generally more robust across different prompts, but are difficult to ‘steer’ in their judgments; that some judge models can be easily fooled by dummy answers such as ‘Yes’ and ‘Sure’; and that judge models are better at detecting completely incorrect answers than partially incorrect ones.
@@
Overall, this work adds to the realm of LLM evaluation research by assessing judges within a clearly defined and objective framework. Our results highlight the utility of using some LLMs as judges but also urge caution in blindly trusting their judgments, even if they are found to be well-aligned with humans. For practitioners using LLMs as judges – regardless of the setup – we recommend not only computing percent agreement, but also Scott’s $\mathbf{\pi}$, and pairing these with a qualitative analysis to ensure that conclusions from judge models are less susceptible to biases. We further elaborate on the limitations of our work in Appendix A. In the future, we plan to expand our work to increasingly more complex scenarios with more open-ended answers and variability and more generally assess how consistent our findings are across dataset samples, benchmarks, and prompt templates.
7 Acknowledgements
We would like to thank Mark Tygert for pointing us to the limitations of Cohen’s Kappa and for helping us proofread Appendix B. AST, KC and VSR received assistance from GPU facilities provided by the University of Massachusetts, Amherst, as well as GPT-4 funding was from the Center of Data Science at Manning College of Information and Computer Science, UMass Amherst.
@@ Appendix A Limitations
For Llama2 7B The average alignment among human evaluators had a Scott’s $\mathbf{\pi}$ of $96.36\pm 1.46$,and the average percent agreement was $98.33\%\pm 0.76\%$. For Llama3.1 70B, we noted that the average alignment among human evaluators had Scott’s $\mathbf{\pi}$ of $95.78\pm 0.30$,and the average percent agreement was $98.72\%\pm 0.10\%$. Given the similarity of these two numbers, we believe that these 1200 samples provide an adequate estimate. In the paper, we take the average.
@@
As each of the nine exam-taker models requires human annotations for each sample, we restricted our analysis to 400 samples in total. This sample size also allowed us to conduct manual annotations and error analysis within 75 human hours/200 GPU hours (see Appendix H) and give reliable confidence intervals while also providing the flexibility to compare a range of models. We were not able to increase the size due to the cost, but a statistical analysis (details provided in Appendix I) illustrated that the variance because of this sample size was very low.
@@ Appendix B A brief explanation of the theoretical issues with Cohen’s kappa
$p_{e}=\sum_{k}\widehat{p_{k12}}=^{ind}\sum_{k}\widehat{p_{k1}}\widehat{p_{k2}}=\sum_{k}\frac{n_{k1}}{N}\frac{n_{k2}}{N}=\frac{1}{N^{2}}\sum_{k}n_{k1}n_{k2},$ (2)
@@ Appendix C Model and dataset details
In Table 3, we show the different models and datasets used in our experiments, along with version and license details.
@@ Appendix D Model evaluation prompt templates
In Figure 7 and Figure 8, we show the prompt templates used for the base and chat exam-taker models during the question answering process.

### v6

@@ 4 Results
In this section we discuss our main results, primarily focusing on the relationship between evaluations by various judge models and human evaluations (§˜4.1), and how that impacts their usability (§˜4.2). To do so, we evaluate their alignment with human judgment and assess how differently they rank the nine exam-taker models compared to humans. In Section 5, we further analyse their precision and recall to further investigate the types of errors that can be made by various judge models. Details about compute requirements and others costs for experiments are given in Appendix˜H.
@@ 4.1 Alignment between judge models and humans
We start by computing Scott’s $\mathbf{\pi}$ scores and percent agreement between the evaluations of each judge model and the human annotators. We show the result in Figure˜1. We observe that percent alignment is high for virtually all models, with the exception of Gemma 2B and EM. Scott’s $\mathbf{\pi}$, on the other hand, has low values for most models, though its value is in the high 80s for Llama-3 70B, Llama-3.1 70B and GPT-4 Turbo. Nevertheless, there still is a significant disparity between human judgment and judge models: the best scoring judge, Llama-3 70B, is 8 points behind human judgment. Notably, EM has the most variance in alignment, while Gemma 2B has the lowest alignment amongst all judges.
@@
To understand how indicative the two alignment metrics are of the expected accuracy of the overall judgement of the models, we plot, for each judge model and exam-taker model, the difference between the score assigned by the judge and the score assigned by a human. In the figure, we can see that for Scott’s $\mathbf{\pi}$ values higher than 80, the evaluation scores are comparatively close to the human evaluation scores, with a difference of up to 5 points in their assigned scores (complete results table provided in Appendix˜J). For percent alignment, on the other hand, even judges that have more than 90% may still differ more than 10 points in their assigned score. Interestingly, the deviation from human-judgements for a single judge model can be quite different depending on the exam-taker model. In Figure˜1(a), Gemma 2B, for instance, sometimes assigns higher scores than humans, and sometimes much lower. In the next section, we further explore this particular pattern.
@@ 4.2 Exploring consistent patterns in judge models
In the previous section, we saw that none of the judge models were as aligned with humans as humans were with each other. As shown in Figure˜2, even the best-aligned judge models can differ by up to 5 points from human-assigned scores. While this limits their ability to perfectly estimate exam-taker model capabilities, judge models can still provide valuable insights to differentiate between exam-taker models. For example, judges with consistent biases may not assign identical scores but could rank models similarly, akin to a very strict teacher.
To assess this, we compare the rankings given by each judge model to the nine exam-taker models, computing Spearman’s rank correlation coefficients $\rho$ (Spearman, 1904) with the human ranking. The rankings are shown in Figure˜3(a), with $\rho$ and $\sigma$ values in Appendix˜L. Most judge models have rank correlations above 0.7, indicating they struggle to distinguish poorer models but do well with better ones. Notably, models like contains and Mistral 7B, which have divergent scores from humans, show high rank correlation ($\rho$ of 0.99 and 0.98, respectively), performing similarly to GPT-4 Turbo and outperforming the better Llama models – though with lower significance values – indicating that identifying which models are better should not be equated to assigning them the correct score.
@@ 5 Analysis
To better understand the judge models, we conduct multiple case studies aimed at identifying common errors and vulnerabilities in the judges we investigate. Specifically, we study their precision and recall and error types (§˜5.1), their sensitivity to the instruction prompt prompt (§˜5.2), how they respond to controlled resposes of specific types (§˜5.3), and the extent to which they have a leniency bias (§˜5.4).
We first examine the precision and recall of the judge models. As shown in Figure˜4(a), both metrics increase moderately with alignment. Figure˜3(b) reveals a similar trend, with a clearer distribution of false positives and negatives. True positives remain consistent across varying judge quality, whereas true negatives exhibit a slight decline as judge quality decreases. Notably, a reduction in judge quality leads to an increase in false positives.
Next, we analyze the errors made by judge models by manually annotating 900 outputs from Llama-7B Base, focusing on top performers GPT-4 Turbo and Llama-3;70B. We categorize error types and determine how often they are correctly judged as incorrect. The results in Table˜2 show that both GPT-4 Turbo and Llama-3;70B excel at identifying answers referring to incorrect entities or containing too many entities. Under-specified and incorrect answers are more challenging, with GPT-4 Turbo performing better on answers with fewer entities than Llama-3;70B.
@@ 5.2 Judge model sensitivity to prompt length and specificity
Next, we investigate how prompt length and specificity affect judge models’ inferences to determine whether their performance is influenced by specificity of the prompt. We use four prompt versions with varying length and specificity.
The first two prompts (Without;guidelines;V1/V2, 45 and 58 tokens) ask for an evaluation without further details. The longer prompts (Guidelines;without;examples and Guidelines;with;examples, 245 and 301 tokens) provide more elaborate guidance and examples. All prompts are listed in Appendix˜M.
Figure˜4(b) shows that GPT-4 Turbo, Llama-3;70B, and Llama-3.1;70B exhibit low variance in human agreement as prompt length and specificity increases. Top performers show high alignment with humans even with minimal instructions, while they slightly improve with more detailed prompts. In contrast, other models lose alignment with increased instructions, likely due to difficulty processing complex instructions.
In a follow-up experiment, we investigate the impact of reference order (see Appendix˜N). Figure˜14 and Figure˜15 shows that larger models maintain consistent judgments regardless of reference order, while smaller models, except Mistral;7B, are more sensitive to it.
@@ 5.3 Evaluating controlled responses
We conduct simple tests on the judge models by having them evaluate dummy benchmark responses. In the first test, the answer is a verbatim reference from the dataset (always correct). In the next three tests, the answers are incorrect. For the second and third tests, the dummy exam-taker model responds with ‘‘Yes’’, and ‘‘Sure’’ respectively. In the fourth test, the evaluated answer is a repetition of the question.
In Figure˜5, we observe that while some judge models correctly identify and mark answers as correct (first test) or incorrect (next three tests), others, like Llama-2;70B, incorrect evaluate many dummy answers, despite showing high human alignment on benchmark evaluations (see Figure˜1(b)). We hypothesize that when the answers are plausible but incorrect, judges can correctly identify them as wrong by comparing them with the reference. However, when the answer is unrelated (e.g., ‘‘Yes’’, and ‘‘Sure’’), judge models may mistakenly mark them as correct, though further research is needed to clarify this behavior.
@@ 5.4 Leniency bias in judge models
Lastly, to get a general sense of the inherent biases or misalignment in the evaluation criteria that might be present in the judge models, we estimate if they have a positive or negative bias in their judgment. To do so, we assume that a judge assigns the correct judgment (i.e. same evaluation as the ground truth) with a probability of $P_{c}$ and assigns the rest of the samples to be ‘‘correct’’ with a probability $P_{+}$, which we call their leniency bias. We estimate the values of $P_{c}$ and $P_{+}$ from the benchmark results444 The theoretical derivation of the expressions for $P_{c}$ and $P_{+}$, as well as the empirical validation for their estimated values can be found in Appendix O. and show them in Figure˜16(a). We observe that $P_{+}$ for most models is significantly higher than $0.5$ (Figure˜16(b)), indicating a tendency of the judge models to evaluate responses as ‘‘correct’’ when their evaluation criteria are not completely aligned with the provided instructions.
@@ 6 Conclusion
In this work, we conduct an extensive study of LLMs as judges, comparing them to human judges and automated evaluation methods. By focusing on a clean evaluation scenario with high inter-human agreement, we identify potential issues with the LLM-as-a-judge paradigm, separate from task ambiguity.
We find that smaller, cost-efficient models, like Mistral;7B, are less effective than larger models such as GPT-4 Turbo, Llama-3.1;70B, and Llama-3;70B, which are better aligned but still fall short of human alignment. Even with high alignment, their scores can differ by up to 5 points from human scores, highlighting the need for caution when using judges in more complex scenarios. We also note that the commonly used metric of percent aligned fails to differentiate between judges effectively. We suggest future work adopt the more robust Scott’s $\mathbf{\pi}$ metric for better distinction.
@@
Next, we note that high alignment scores are not always necessary to discriminate between models. While GPT-4 Turbo and Llama-3 have excellent alignment scores, simpler and more cost-efficient models, like contains, perform similarly in ranking exam-taker models, despite lower alignment scores and score deviations. For studies focused on ranking models rather than estimating exact scores, these approaches can be as suitable as more expensive ones.
@@
Lastly, we run experiments to assess judge models’ sensitivity to prompts, precision, recall, error types, leniency, and vulnerability to dummy answers. We find that smaller models are more likely to judge positively when in doubt, that lower-alignment models lack precision, and that better models are more robust across different prompts but harder to "steer." Some judge models are easily fooled by dummy answers like ’’Yes’’ and ’’Sure’’ and are better at detecting completely incorrect answers than partially incorrect ones.
@@
Overall, this work contributes to LLM evaluation by assessing judges in a clearly defined framework. Our results highlight the potential of LLMs as judges but caution against blindly trusting their judgments, even when aligned with humans. We recommend computing both percent agreement and Scott’s $\mathbf{\pi}$, paired with qualitative analysis, to avoid bias. We discuss limitations in Appendix˜A and plan to expand our work to more complex scenarios in the future.
@@ Appendix A Limitations
For Llama2 7B The average alignment among human evaluators had a Scott’s $\mathbf{\pi}$ of $96.36\pm 1.46$,and the average percent agreement was $98.33\%\pm 0.76\%$. For Llama3.1 70B, we noted that the average alignment among human evaluators had Scott’s $\mathbf{\pi}$ of $95.78\pm 0.30$,% and the average percent agreement was $98.72\%\pm 0.10\%$. Given the similarity of these two numbers, we believe that these 1200 samples provide an adequate estimate. In the paper, we take the average.
@@
As each of the nine exam-taker models requires human annotations for each sample, we restricted our analysis to 400 samples in total. This sample size also allowed us to conduct manual annotations and error analysis within 75 human hours/200 GPU hours (see Appendix˜H) and give reliable confidence intervals while also providing the flexibility to compare a range of models. We were not able to increase the size due to the cost, but a statistical analysis (details provided in Appendix˜I) illustrated that the variance because of this sample size was very low.
@@ Appendix B A brief explanation of the theoretical issues with Cohen’s kappa
$p_{e}$ $=\sum_{k}\widehat{p_{k12}}=^{ind}\sum_{k}\widehat{p_{k1}}\widehat{p_{k2}}$
$=\sum_{k}\frac{n_{k1}}{N}\cdot\frac{n_{k2}}{N}=\frac{1}{N^{2}}\sum_{k}n_{k1}n_{k2}$
@@ Appendix C Model and dataset details
In Table˜3, we show the different models and datasets used in our experiments, along with version and license details.
@@ Appendix D Model evaluation prompt templates
In Figure˜6 and Figure˜7, we show the prompt templates used for the base and chat exam-taker models during the question answering process.